A nut mixture is made with almonds that cost $8.00 per pound and peanuts that cost $3.00 per pound. The mixture costs $5.00 per pound. How many pounds of each type of nut should be used to make 20 pounds of this mixture?
Answer by jorel1380(3719)

m+n=20
8m+3n=20(5)=100
3m+3n=60
5m=40
m=8
n=12
You need 8 lbs at $8/lb. and 12 lbs at $3/lb. to make 20 lbs at $5/lb..